OpenStudy (anonymous):

(1/x)-(4/3x)=(1/x-2) need to solve equation

OpenStudy (lgbasallote):

\[\Large \frac 1x - \frac{4}{3x} = \frac{1}{x-2}\] is this the question?

OpenStudy (anonymous):

nope

OpenStudy (lgbasallote):

\[\Large \frac 1x - \frac{4}{3x} = \frac 1x - 2\]

OpenStudy (anonymous):

yes it is soory

OpenStudy (anonymous):

sorry

OpenStudy (lgbasallote):

lol you confuse me

OpenStudy (lgbasallote):

so it's the first one?

OpenStudy (anonymous):

well the last part is (1/x-2)

OpenStudy (anonymous):

the x-2 is the denominator

OpenStudy (lgbasallote):

so is it first one or second one?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

first

OpenStudy (lgbasallote):

start by converting the left side to common denominators \[\frac 1x = \frac{3}{3x}\] therefore i will turn the equation inot \[\frac{3}{3x} - \frac{4}{3x} = \frac{1}{x-2}\] \[\frac{3-4}{3x} = \frac{1}{x-2}\] \[\frac{-1}{3x} = \frac{1}{x-2}\] cross multiply \[-(x-2) = 3x\] \[-x + 2 = 3x\] add x to both sides \[2 = 3x +x\] \[2 = 4x\] can you solve for x now?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

great